# MaTe3D: Mask-guided Text-based 3D-aware Portrait Editing

### v3

@@ V-A2 Setting of Editing
FENeRF [10] 29.0 3.728 0.61 / /
StyleNeRF [7] 7.8 0.220 0.62 14.9 0.368
StyleSDF [6] 13.1 0.269 0.74 12.8 0.455
MVCGAN [40] 13.4 0.375 0.58 17.1 0.198
EG3D [5] 4.7 0.132 0.77 2.7 0.041
IDE-3D [11] 4.6 0.130 0.76 / /
BallGAN [41] 5.7 0.312 0.75 4.7 0.096
MaTe3D(w/o dual tri-planes) 5.8 0.389 0.75 3.9 0.106
@@
TABLE I: Quantitative comparison with the state-of-the-art methods. The comparison results are quoted from IDE-3D [11], EG3D [5] and BallGAN [41].
V-B CatMask-HQ
Current mask-based editing methods are primarily validated only on CelebA-HQ Mask [42] or FFHQ [1]While some few-shot techniques have produced faces and masks in other domains, such as cat faces [43], both the image and the mask quality are subpar with noticeable aliasing [44].
@@ V-B CatMask-HQ
IDE-3D [11] w nose 3039.3770 0.0443
IDE-3D [11] w nose, mouth 3084.0366 0.0445
IDE-3D [11] w nose, mouth, hair 2607.2324 0.0450
IDE-3D [11] w nose, mouth, hair, ear 2677.9805 0.0451
@@
TABLE II: Evaluation on Chamfer-distance (L1) and normal consistency score (Norm.) [45] between representative editable 3D-aware generative model with MaTe3D.
V-C Baselines
We evaluate the generator of MaTe3D against state-of-the-art genertors, including FENeRF [10], StyleNeRF [7], StyleSDF [6], EG3D [5], MVCGAN [40], BallGAN [41] and IDE-3D [11], respectively. We evaluate MaTe3D from two aspects including text-based and mask-text-hybrid editing. For text-based editing, we compare MaTe3D with LENeRF, FENeRF+StyleCLIP, IDE-3D+StyleCLIP, FaceDNeRF [24] and HeadSculpt [32], respectively. We reimplement LENeRF [12] and HeadSculpt [32] strictly according to their papers. We reproduce FENeRF+StyleCLIP and IDE-3D+StyleCLIP for comparison purposes, similarly to LENeRF [12]. In hybrid editing, we compare two baselines: FENeRF+StyleCLIP and IDE-3D+StyleCLIP with mask and text editing simultaneously.
V-D MaTe3D generator Evaluation
@@
We adopt three metrics to comprehensively evaluate the synthesis quality of generated images quantitatively. Both the lower Frechet Inception Distance (FID) [46] and Kernel Inception Distance (KID) [47] indicate higher quality of synthesized images. Multi-view identity consistency (ID) [48] calculates the face similarity under different sampled camera pose. As shown in Tab. I, MaTe3D achieves comparable results with SOTAs in terms of FID, KID and ID. Fig. 3 presents a qualitative comparison of MaTe3D against 3D-aware GANs. It is evident that MaTe3D achieves superior image quality and better view consistency than the baselines, indicating the effectiveness of both the global and the local SDF representations in the generator. Also, we present additional results for CatMask-HQ, and compare the image quality with that by EG3D [5], as depicted in Fig. 3 and Tab. I. Our method produces better image quality with better view consistency.
@@
Decompositional Geometries: qualitative result.   We compare MaTe3D with two other state-of-the-art methods: FENeRF and IDE-3D. As shown in Fig. 4, MaTe3D outperforms its predecessors. FENeRF produces visually irritating staircasing artifacts [49] (the so-called bulls eye effect [50]) that are particularly noticeable when viewing small objects up close. It is possibly caused by the sparse and clustered sampling points in the frustum of FENeRF. IDE-3D produces photorealistic semantic-aware images of relatively compact whole faces. However, it fails to maintain a well-established zero level-set in its neural radiance field, resulting in constant density distribution along rays even for local parts like nose, mouth and eyes. The situation may be caused by ignoring the association between geometries and semantic information, which leads to the mis-disentanglement of underlying local geometries.
Decompositional Geometries: quantitative result.   We aim to investigate the compositional consistency in Mate3D in quantitative way. We combine the local facial meshes into a unified mesh, and then compare it with the original whole face. The results, presented in Tab. II, demonstrate that MaTe3D significantly enhances 3D surface reconstruction quality compared to IDE-3D. Note that we do not include FENeRF in this comparison as its produced geometries have different resolutions of MaTe3D and IDE-3D, making it an unfair comparison. MaTe3D outperforms IDE-3D by large margins in terms of Chamfer-L1 (approximately 1/5 of IDE-3D) and normal consistency (approximately two times of IDE-3D), indicating its superior performance in integrity and compositionality.
V-E Edit Results
@@
TABLE III: Quantitative comparison with the state-of-the-art editing methods. SC in table represents StyleCLIP.
@@
We report the CLIPScore [51] values in Tab. III, which replies the correlation between edited images and input texts. We achieve the best ClIPscore ($0.791$) comparing with LENeRF ($0.612$), FENeRF+StyleCLIP ($0.729$), and IDE-3D+StyleCLIP ($0.762$), respectively.
@@
Figure 10: Ablation study on the utilization of Condition Distillation on Geometry and Texture (CDGT). The absence of Condition Updating Scheme in CDGT results in unstable texture and noticeable blurring during editing. Without combining gradients or naively mixing them with the TADA [38] strategy, the geometry remains unconvincing and hollow, leading to a degradation in 3D quality. Our method demonstrates stable textures and convincing geometry.
@@
V-F Ablation Study
Local SDF Representations.   We conduct an ablation study on the local SDF representation. As depicted in Fig. 4, Our model (w/o SDFs) generates photorealistic images and semantic-aware masks, but suffers from unsatisfactory geometries. Benefiting from our generator architecture, the results without SDFs have superior geometries compared to IDE-3D [11], but errors still exist in modeling eyeglasses (incomplete edges) and nose (longer and deeper than normal). These errors occur due to a loss of connection between decompositional facial components. The poor geometry also leads to inaccurate masks.
@@ V-F Ablation Study
Consistency Loss and Network Architecture.   As depicted in Fig. 9, we perform an ablation study on the proposed consistency loss and network architecture. Our findings suggest that omitting the density consistency loss leads to a significant degradation in image quality, severely compromising the overall fidelity of the whole image. Additionally, omitting the SDF consistency loss results in aliasing edges in the semantic mask and requires more iterations to converge. Incorporating dual tri-plane generation improves image quality (see Tab. I) and provides more details in specific areas, such as glasses in face (see Fig. 9).
Conditional Distillation on Geometry and Texture.   As shown in Fig. 10, we conduct an ablation study on our proposed blending Conditional Distillation on Geometry and Texture (CDGT). Our observations indicate that without Condition Updating Scheme in CDGT, the editing phase loses 3D-aware control, leading to unstable textures, and noticeable blurring. Without gradient combination, image synthesis in novel views results in severe artifacts due to 3D degradation caused by mismatch between the underlying geometry and texture. The strategy proposed in TADA [38] can partially alleviate this issue. However, the results still suffer from unstable texture and lead to hollow geometric structures. By incorporating gradient combination via random blend, we introduce randomness into the fusion process. This random blending boosts the resilience of edited results by preventing overfitting on specific patterns while enhancing its capability to handle challenging discrete attributes such as hats and glasses.
@@
VI Applications
VI-A Real Portrait Editing
Our MaTe3D allows for high-fidelity 3D inversion and mask-guided text-based editing on real portrait images. We start by projecting the input image into the latent space using PTI [52], and then use MaTe3D for mask-guided and text-based editing. The results are displayed in Fig. 11, demonstrating that our approach can achieve high-quality 3D portrait editing on real images.
VI-B Out-of-Domain Editing
@@
VI-C Face Swapping
Face swapping is a crucial research area in computer vision with extensive applications in entertainment industry [53]. Our method can also perform face swap editing with celebrities. As shown in Fig. 13, with a prompt of a celebrity, MaTe3D can transfer their face to the source image while keeping the remaining regions unchanged.
VII Conclusion

### v4

@@ 5.1.2 Setting of Editing
FENeRF fenerf 29.0 3.728 0.61 / /
StyleNeRF stylenerf 7.8 0.220 0.62 14.9 0.368
StyleSDF stylesdf 13.1 0.269 0.74 12.8 0.455
MVCGAN mvcgan 13.4 0.375 0.58 17.1 0.198
EG3D eg3d 4.7 0.132 0.77 2.7 0.041
IDE-3D ide3d 4.6 0.130 0.76 / /
BallGAN ballgan 5.7 0.312 0.75 4.7 0.096
MaTe3D
(w/o dual tri-planes) 5.8 0.389 0.75 3.9 0.106
@@
Table 1: Quantitative comparison with the state-of-the-art methods. The comparison results are quoted from IDE-3D ide3d , EG3D eg3d and BallGAN ballgan .
5.2 CatMask-HQ
Current mask-based editing methods are primarily validated only on CelebA-HQ Mask maskgan or FFHQ stylegan While some few-shot techniques have produced faces and masks in other domains, such as cat faces sem2nerf , both the image and the mask quality are subpar with noticeable aliasing kangle2023pix2pix3d .
@@ 5.2 CatMask-HQ
IDE-3D ide3d w nose 3039.3770 0.0443
IDE-3D ide3d w nose, mouth 3084.0366 0.0445
IDE-3D ide3d w nose, mouth, hair 2607.2324 0.0450
IDE-3D ide3d w nose, mouth, hair, ear 2677.9805 0.0451
@@
Table 2: Evaluation on Chamfer-distance (L1) and normal consistency score (Norm.) mescheder2019occupancy between representative editable 3D-aware generative model with MaTe3D.
5.3 Baselines
We evaluate the generator of MaTe3D against state-of-the-art genertors, including FENeRF fenerf , StyleNeRF stylenerf , StyleSDF stylesdf , EG3D eg3d , MVCGAN mvcgan , BallGAN ballgan and IDE-3D ide3d , respectively. We evaluate MaTe3D from two aspects including text-based and mask-text-hybrid editing. For text-based editing, we compare MaTe3D with LENeRF, FENeRF+StyleCLIP, IDE-3D+StyleCLIP, FaceDNeRF facednerf and HeadSculpt headsculpt , respectively. We reimplement LENeRF lenerf and HeadSculpt headsculpt strictly according to their papers. We reproduce FENeRF+StyleCLIP and IDE-3D+StyleCLIP for comparison purposes, similarly to LENeRF lenerf . In hybrid editing, we compare two baselines: FENeRF+StyleCLIP and IDE-3D+StyleCLIP with mask and text editing simultaneously.
5.4 MaTe3D generator Evaluation
@@
We adopt three metrics to comprehensively evaluate the synthesis quality of generated images quantitatively. Both the lower Frechet Inception Distance (FID) fid and Kernel Inception Distance (KID) kid indicate higher quality of synthesized images. Multi-view identity consistency (ID) id calculates the face similarity under different sampled camera pose. As shown in Tab. 1, MaTe3D achieves comparable results with SOTAs in terms of FID, KID and ID. Fig. 3 presents a qualitative comparison of MaTe3D against 3D-aware GANs. It is evident that MaTe3D achieves superior image quality and better view consistency than the baselines, indicating the effectiveness of both the global and the local SDF representations in the generator. Also, we present additional results for CatMask-HQ, and compare the image quality with that by EG3D eg3d , as depicted in Fig. 3 and Tab. 1. Our method produces better image quality with better view consistency.
@@
Decompositional Geometries: qualitative result.   We compare MaTe3D with two other state-of-the-art methods: FENeRF and IDE-3D. As shown in Fig. 4, MaTe3D outperforms its predecessors. FENeRF produces visually irritating staircasing artifacts morsy2022shape (the so-called bulls eye effect IDW1968 ) that are particularly noticeable when viewing small objects up close. It is possibly caused by the sparse and clustered sampling points in the frustum of FENeRF. IDE-3D produces photorealistic semantic-aware images of relatively compact whole faces. However, it fails to maintain a well-established zero level-set in its neural radiance field, resulting in constant density distribution along rays even for local parts like nose, mouth and eyes. The situation may be caused by ignoring the association between geometries and semantic information, which leads to the mis-disentanglement of underlying local geometries.
Decompositional Geometries: quantitative result.   We aim to investigate the compositional consistency in Mate3D in quantitative way. We combine the local facial meshes into a unified mesh, and then compare it with the original whole face. The results, presented in Tab. 2, demonstrate that MaTe3D significantly enhances 3D surface reconstruction quality compared to IDE-3D. Note that we do not include FENeRF in this comparison as its produced geometries have different resolutions of MaTe3D and IDE-3D, making it an unfair comparison. MaTe3D outperforms IDE-3D by large margins in terms of Chamfer-L1 (approximately 1/5 of IDE-3D) and normal consistency (approximately two times of IDE-3D), indicating its superior performance in integrity and compositionality.
5.5 Edit Results
@@
Table 3: Quantitative comparison with the state-of-the-art editing methods. SC in table represents StyleCLIP.
@@
We report the CLIPScore clipscore values in Tab. 3, which replies the correlation between edited images and input texts. We achieve the best ClIPscore ($0.791$) comparing with LENeRF ($0.612$), FENeRF+StyleCLIP ($0.729$), and IDE-3D+StyleCLIP ($0.762$), respectively.
@@
Figure 10: Ablation study on the utilization of Condition Distillation on Geometry and Texture (CDGT). The absence of Condition Updating Scheme in CDGT results in unstable texture and noticeable blurring during editing. Without combining gradients or naively mixing them with the TADA tada strategy, the geometry remains unconvincing and hollow, leading to a degradation in 3D quality. Our method demonstrates stable textures and convincing geometry.
@@
5.6 Ablation Study
Local SDF Representations.   We conduct an ablation study on the local SDF representation. As depicted in Fig. 4, Our model (w/o SDFs) generates photorealistic images and semantic-aware masks, but suffers from unsatisfactory geometries. Benefiting from our generator architecture, the results without SDFs have superior geometries compared to IDE-3D ide3d , but errors still exist in modeling eyeglasses (incomplete edges) and nose (longer and deeper than normal). These errors occur due to a loss of connection between decompositional facial components. The poor geometry also leads to inaccurate masks.
@@ 5.6 Ablation Study
Consistency Loss and Network Architecture.   As depicted in Fig. 9, we perform an ablation study on the proposed consistency loss and network architecture. Our findings suggest that omitting the density consistency loss leads to a significant degradation in image quality, severely compromising the overall fidelity of the whole image. Additionally, omitting the SDF consistency loss results in aliasing edges in the semantic mask and requires more iterations to converge. Incorporating dual tri-plane generation improves image quality (see Tab. 1) and provides more details in specific areas, such as glasses in face (see Fig. 9).
Conditional Distillation on Geometry and Texture.   As shown in Fig. 10, we conduct an ablation study on our proposed blending Conditional Distillation on Geometry and Texture (CDGT). Our observations indicate that without Condition Updating Scheme in CDGT, the editing phase loses 3D-aware control, leading to unstable textures, and noticeable blurring. Without gradient combination, image synthesis in novel views results in severe artifacts due to 3D degradation caused by mismatch between the underlying geometry and texture. The strategy proposed in TADA tada can partially alleviate this issue. However, the results still suffer from unstable texture and lead to hollow geometric structures. By incorporating gradient combination via random blend, we introduce randomness into the fusion process. This random blending boosts the resilience of edited results by preventing overfitting on specific patterns while enhancing its capability to handle challenging discrete attributes such as hats and glasses.
@@
6 Applications
6.1 Real Portrait Editing
Our MaTe3D allows for high-fidelity 3D inversion and mask-guided text-based editing on real portrait images. We start by projecting the input image into the latent space using PTI pti , and then use MaTe3D for mask-guided and text-based editing. The results are displayed in Fig. 11, demonstrating that our approach can achieve high-quality 3D portrait editing on real images.
6.2 Out-of-Domain Editing
@@
6.3 Face Swapping
Face swapping is a crucial research area in computer vision with extensive applications in entertainment industry 3dswap . Our method can also perform face swap editing with celebrities. As shown in Fig. 13, with a prompt of a celebrity, MaTe3D can transfer their face to the source image while keeping the remaining regions unchanged.
7 Conclusion